OpenStudy (anonymous):

Determine if the following is true or false a) fog=gof . b)If f(X) is even function ,then f(x-a) is also even function. please.....

OpenStudy (anonymous):

false

OpenStudy (anonymous):

both are false in general would you like examples?

OpenStudy (anonymous):

for example \[f(x)=x^2,g(x)=x-2\] then \[f\circ g = f(g(x))=f(x-2)=(x-2)^2=x^2-4x+4\] \[g\circ f = g(f(x))=g(x^2)=x^2-2\]

OpenStudy (anonymous):

f(x)=sqrt(x) and g(x)=x^2 can be an example for a.

OpenStudy (anonymous):

for b) we can also use \[f(x)=x^2\] which is even , but \[f(x-2)=x^2-4x+4\] is neither even or odd

OpenStudy (anonymous):

i see totally got it thanks a lot